# Multiple Invertible and Partial-Equivariant Function for Latent Vector Transformation to Enhance Disentanglement in VAEs

### v1

@@ Abstract
Disentanglement learning is a core issue for understanding and re-using trained information in Variational AutoEncoder (VAE), and effective inductive bias has been reported as a key factor. However, the actual implementation of such bias is still vague. In this paper, we propose a novel method, called Multiple Invertible and partial-equivariant transformation (MIPE-transformation), to inject inductive bias by 1) guaranteeing the invertibility of latent-to-latent vector transformation while preserving a certain portion of equivariance of input-to-latent vector transformation, called Invertible and partial-equivariant transformation (IPE-transformation), 2) extending the form of prior and posterior in VAE frameworks to an unrestricted form through a learnable conversion to an approximated exponential family, called Exponential Family conversion (EF-conversion), and 3) integrating multiple units of IPE-transformation and EF-conversion, and their training. In experiments on 3D Cars, 3D Shapes, and dSprites datasets, MIPE-transformation improves the disentanglement performance of state-of-the-art VAEs.
@@ 1 Introduction
Disentanglement learning to learn more interpretable representations is broadly useful in artificial intelligence fields such as classification Singla et al. (2021), zero-shot learning Tenenbaum (2018), and domain adaptation Li et al. (2019); Zou et al. (2020). The disentangled representation is defined as a change in a single dimension, which corresponds to unique semantic information. Several works have been conducted based on this framework.
A major model for enhancing the disentanglement learning is Variational AutoEncoder (VAE) Kingma & Welling (2013). Based on VAE, unsupervised disentangled representation learning has been elaborated Higgins et al. (2017); Chen et al. (2018); Kim & Mnih (2018); Jeong & Song (2019); Li et al. (2020) through the factorizable variations and control of uncorrelatedness of each dimension of representations. Moreover, VAE models that handle the shape of prior as a Gaussian mixture Dilokthanakul et al. (2016b) or von Mises-Fisher Davidson et al. (2018) were also developed, but the disentanglement is still incomplete. As a critical point, there is a report that unsupervised disentanglement learning is impossible without inductive bias Locatello et al. (2019).
Recently, such inductive bias has been introduced in various perspectives on transformation of latent vector space. Intel-VAE Miao et al. (2022) proposed the benefit of invertible transformation of the space to another latent space to provide better data representation, which includes hierarchical representations. Group theory based bias also shows significant improvement on disentanglement Zhu et al. (2021); Yang et al. (2022), whose definition follows Higgins et al. (2018a), which is based on the group theory. The works show that equivariant transformation between input and latent vector space has a key role of disentanglement.
Inspired by the above works, we propose a Multiple Invertible and partial-equivariant transformation (MIPE-transformation) method, which is simply insertable to VAEs. The partial-equivariance is defined by Romero & Lohit (2022) and we follow its definition: $f(g\cdot x)=g\cdot f(x)$ $\forall x\in\mathcal{X},\forall g\in G^{\prime},G^{\prime}\subset G$, where $G^{\prime}$ is a subset of group $G$ ($f$ is partially equivariant to group $G$). First, we assume that an encoder is partial-equivariant and we call it an encoder equivariance condition. The method adopts the matrix exponential to hold the invertible property of latent-to-latent (L2L) vector transformation. Then, we constrain the L2L transformation to a symmetric matrix exponential to be partial-equivariant to a subgroup between latent and transformed latent space. Because it extends the encoder to be partial-equivariant to a subgroup between input and transformed latent space. The IPE-transformation generates an uncertain form of latent vector distributions, so we provide a training procedure to force them to be close to an exponential family, called exponential family conversion (EF-conversion). This conversion enables the uncertain distribution to work in the typical training framework of VAEs. Then, we mathematically show that the multiple uses of IPE-transformation work as $\beta$ parameters Higgins et al. (2017) controlled for enhancing disentanglement learning. Also, we propose the implicit semantic mask to induce a semantic mask in the latent vector space, different to Yang et al. (2020). In experiments with quantitative and qualitative analysis, MIPE-transformation shows significant improvement in disentangled representation learning in 3D Cars, 3D Shapes, and dSprites tasks. Our main contributions are summarized as follows.
@@
    We propose to use a symmetric matrix exponential as a latent-to-latent vector transformation function for inducing inductive bias based on invertible and equivariant properties with mathematical analysis.
@@
    We provide a training procedure and losses for VAEs to learn unknown latent vector distribution as an approximated exponential family.
@@
    We propose the novel MIPE-transformation architecture to integrate multiple IPE-transformation and EF-conversion, which is widely applicable to state-of-the-art VAEs.
@@
    We empirically analyze the properties of MIPE-transformation and validate its effectiveness in disentanglement learning on benchmarks.
Table 1: Terms and Notations
${\bm{z}}$ Latent vector from encoder $\psi(\cdot)$ Invertible function
$\hat{{\bm{z}}}_{m}$ Transformed latent vector by $\psi_{m}(\cdot)$ $\hat{\bm{\epsilon}}_{m}$ Transformed prior samples by $\psi_{m}(\cdot)$
$\bm{\theta}_{\hat{{\bm{z}}}_{m}}$ Natural Parameter of posterior $\bm{\theta}_{\hat{\bm{\epsilon}}_{m}}$ Natural Parameter of prior
$T$ Sufficient Statistics $A$ Log-Normalizer
$\bm{\nu}$ Evidence $D_{\mathrm{KL}}(\cdot||\cdot)$ Kullback-Leibler divergence
$f_{{\bm{x}}}(\cdot)$ Power Density Function $M_{n}(\mathbb{R})$ A set of $n\times n$ real matrix
$GL_{n}(\mathbb{R})$ General Linear Group $Sym_{n}(\mathbb{R})$ A set of $n\times n$ symmetric real matrix
$E_{M}$$\{\mathbf{e}^{M}|M\in M_{n}(\mathbb{R})\}$ $E_{S}$$\{\mathbf{e}^{S}|S\in Sym_{n}(\mathbb{R})\}$
$G_{S}$ $G_{S}:(\mathbf{e}^{S},\ast)$ $G_{I}$ Group of input space for symmetries
$G_{L}$ Group of latent space for symmetries $J$ $G_{S}\cap G_{L}$
$\psi_{M}(\cdot)$ $\psi_{M}(\cdot)\in M_{n}(\mathbb{R})$ $\psi_{E_{M}}(\cdot)$ $\psi_{E_{M}}(\cdot)\in E_{M}$
$\psi_{E_{S}}(\cdot)$ $\psi_{E_{S}}(\cdot)\in E_{S}$ $\mathbf{0}$ zero vector
$\mathbf{0}_{n,n}$ n by n zero matrix $\mathcal{X}$ Input space
$\mathcal{Z}$ Latent vector space $\hat{\mathcal{Z}}$ Transformed latent vector space
$\Xi$ $G_{I}\times G_{L}\rightarrow G_{L}$ $\Gamma$ $G_{L}\times G_{T}\rightarrow G_{T}$
$\Xi^{J}$ $G_{I}^{J}\times G_{L}^{J}\rightarrow G_{L}^{J}$ $\Gamma^{J}$ $G_{L}^{J}\times G_{T}^{J}\rightarrow G_{T}^{J}$
@@ 2 Related Work
Recently, various works have focused on the unsupervised disentanglement learning. Previous works are based on Bengio et al. (2013) definition. One of the branches is InfoGAN Chen et al. (2016) based works such as IB-GAN Jeon et al. (2021) implement extra regularizer to improve informativeness Eastwood & Williams (2018). The other branch is based on the VAE. $\beta$-VAE Higgins et al. (2017) penalizes Kullback-Leibler divergence (KL divergence) with weighted hyper-parameters. Factor VAE Kim & Mnih (2018) and $\beta$-TCVAE Chen et al. (2018) are trained with total correlation (TC) to make independent dimensions on a latent vector with discriminator and divided components of KL divergence term. Differently, we consider the recent disentanglement definition based on group theory Higgins et al. (2018a).
Following the definitions of disentangled representation learning by group theory, several works have emphasized equivariant and improved disentangled representation learning. Commutative Lie Group VAE (CLG-VAE) Zhu et al. (2021) proposed direct mapping of the latent vector into Lie algebra to obtain group structure (inductive bias) with constraints: commutative and hessian loss. Furthermore, Groupified VAE Yang et al. (2022) utilizes Spatial Broadcast Decoder Watters et al. (2019) to implement an equivariant function to the cyclic group with guaranteeing commutativity and invertibility of group actions. Topographic VAE Keller & Welling (2021) combines Student’s-t distributions and variational inference. It enforces rotated latent vectors to be equivariant. On the other hand, we apply unrestricted prior and posterior for disentanglement learning.
There are several inductive biases to learning unsupervised disentanglement, such as group theory based and sequential order. In this section, we briefly discuss sequential order inductive bias even though its method is considered in different domains such as text and video frames. To individualize the static (time-invariant) and dynamic (time-variant), Li & Mandt (2018); Bai et al. (2021) proposed the latent variables one ($f$) is only dependent on the given times series datasets $x_{1:T}$, and the other ($\mathbf{z}_{1:T})$ is dependent on the $x_{1:T}$ and $f$. Moreover Bai et al. (2021) propose the novel ELBO with maximizing mutual information between the input and the latent vectors. These works empirically show that sequential order which includes separated latent vectors improves unsupervised disentanglement learning with diverse qualitative analysis. Differently in group theory based approaches, the proposed methods consider equivariant function between input and latent vector space.
@@
Other VAE approaches implement other prior from Gaussian distribution to transformed Gaussian distribution, Gaussian mixture distribution Dilokthanakul et al. (2016a) or von Mises-Fisher distribution Davidson et al. (2018). InteL-VAE Miao et al. (2022) shows that transformed Gaussian distribution by the invertible function trains hierarchical representation with manual function. We show more clear relation of invertibility to disentanglement and improve VAEs to use its unrestricted form of prior.
Invertible and equivariant Deep Neural Networks have been investigated with normalizing flows. As proven by Xiao & Liu (2020), utilized matrix exponential on Neural networks is invertible, but it only provides mathematical foundations of the tra nsformation. Matrix exponential is utilized to implement an invertible and equivariant function to improve the generative flow compare to linear function Hoogeboom et al. (2020). To specify the exponential familyt, other works contribute uncertainty of exponential family distribution with Bayesian update Charpentier et al. (2020; 2022b). In addition, Ranganath et al. (2014) hierarchically controls the natural parameter across the layers and determines the exponential family distribution with the moment of sufficient statistic. In our work, we show how to use it for disentanglement learning.
Figure 1: The overall architecture of our proposed MIPET-VAE. The invertible and partial-equivariant function $\psi(\cdot)$ for L2L transformation consists of a symmetric matrix exponential to be 1) invertible and 2) partial-equivariant. Then 3) EF conversion module converges the distribution of unrestricted $\hat{{\bm{z}}}$ to be EF with $\mathcal{L}_{el}$ loss. Also, it applies KL divergence loss ($\mathcal{L}_{kl}$) between the transformed posterior and prior, which are expressed by the power density function of EF. In the last, EF conversion reduces the computational error ($\mathcal{L}_{cali}$) between approximated and true KL divergence. 4) The reddish color represents the integration parts. The blue figures represent each property. The details of the gray box are in Figure 2.
@@ 3 Method
The overview of a VAE equipped with MIPE-transformation is shown in Figure 1. The MIPE-transformation has three main components: 1) IPE-transformation Unit to transform latent vectors with invertible and partial-equivariant properties, 2) EF-conversion Unit to extend VAEs to learn the exponential family distribution of latent vectors, and 3) integrated training and generation process for multiple uses of IPE-transformation and EF-conversion.
3.1 Invertible and Partial-Equivariant Function for L2L Transformation
3.1.1 Invertible Property by Using Matrix Exponential
To guarantee the invertible property of IPE-transformation, we use a function $\psi(\cdot)=\mathbf{e}^{{\bm{M}}}\ast\cdot$ for the transformation, where ${\bm{M}}$ is in $n\times n$ real number matrix set $M_{n}(\mathbb{R})$ Xiao & Liu (2020). The operator $\ast$ is matrix multiplication, and $\mathbf{e}^{{\bm{M}}}=\sum_{k}^{\infty}\frac{{\bm{M}}^{k}}{k!}$. Our motivation is to use the benefits of injecting explicit inductive bias for disentanglement Locatello et al. (2019); Miao et al. (2022). InteL-VAE effectively extracts hierarchical representation, which includes low-level features (affect to a specific factor) and high-level features (affect to complex factors) with an invertible transformation function Miao et al. (2022).
Figure 2: $G_{I}$, and $G_{L}$ are obtained through encoder $q_{\phi}$ (encoder equivariance condition). The left side figure shows the relation between each space and symmetries. If $\psi(\cdot)$ is equivariant function over all $G_{L}$, and $G_{T}$, then there exist $\Gamma$, where $\Gamma:G_{L}\rightarrow G_{T}$, and $\Xi\circ\Gamma:G_{I}\rightarrow G_{T}$. However, unrestricted $\psi(\cdot)$ has no guarantee to be partial- or full-equivariant. The red arrows represent our method: L2L transformation guarantees $\Gamma^{J}:G_{L}^{J}\rightarrow G_{T}^{J}$, and $\Xi^{J}\circ\Gamma^{J}:G_{I}^{J}\rightarrow G_{T}^{J}$, given the encoder equivariance condition $\Xi:G_{I}\rightarrow G_{L}$.
3.1.2 Why Should L2L Transformation Be Equivariant?
Let’s consider equivariant function between the input and transformed latent vector space, directly used for a decoder in the VAE frameworks. All L2L transformations do not extend the encoder equivariance condition to the relation between input and transformed latent space. This problem is more precisely shown in Figure 2, which illustrates partial equivariance condition over the input space $\mathcal{X}$, latent vector space $\mathcal{Z}$, and its transformed latent vector space $\hat{\mathcal{Z}}$ with a corresponding group of symmetries $G_{I}$, $G_{L}$, and $G_{T}$, respectively. In the VAEs literature, it has not been reported to restrict L2L transformation to guarantee equivariant function between two spaces, so we propose a solution to guarantee at least a part of symmetries to be equivariant.
3.1.3 Equivariance Property with Symmetric Matrix Exponential

### v3

@@ Abstract
Disentanglement learning is central to understanding and reusing learned representations in variational autoencoders (VAEs). Although equivariance has been explored in this context, effectively exploiting it for disentanglement remains challenging. In this paper, we propose a novel method, called Multiple Invertible and Partial-Equivariant Transformation (MIPE-Transformation), which integrates two main parts: (1) Invertible and Partial-Equivariant Transformation (IPE-Transformation), guaranteeing an invertible latent-to–transformed-latent mapping while preserving partial input-to-latent equivariance in the transformed latent space; and (2) Exponential-Family Conversion (EF-Conversion) to extend the standard Gaussian prior to an approximate exponential family via a learnable conversion. In experiments on the 3D Cars, 3D Shapes, and dSprites datasets, MIPE-Transformation improves the disentanglement performance of state-of-the-art VAEs.
@@ 1 Introduction
Disentanglement learning aims to learn interpretable representations and has utility in tasks such as classification (Singla et al., 2021), zero-shot learning (Tenenbaum, 2018), and domain adaptation (Li et al., 2019; Zou et al., 2020). A disentangled representation varies along one latent coordinate in a way that corresponds to a single factor of variation while keeping other factors essentially fixed. Among existing approaches, the Variational Autoencoder (VAE) (Kingma and Welling, 2013) is a primary framework; unsupervised methods encourage factorized latents and reduce inter-dimension correlation (Higgins et al., 2017; Chen et al., 2018; Kim and Mnih, 2018; Jeong and Song, 2019; Li et al., 2020), but Locatello et al. (2019) shows that disentanglement is not identifiable without inductive bias.
To inject such bias, prior work follows two lines: 1) statistics-driven methods that replace the standard Gaussian prior with richer, often non-Gaussian alternatives. For example, Kim et al. (2019); Mathieu et al. (2018) investigate Gaussian-mixture and Student-$t$ priors and demonstrate limitations of the standard Gaussian prior for disentanglement. Similarly, Locatello et al. (2020); Togo et al. (2025) consider priors beyond the Gaussian family and report improved disentanglement performance. However, these approaches are grounded in the original notion of disentangled representation (Bengio et al., 2013) rather than the group-theoretic formulation, and they do not explicitly encode equivariance.
The second line comprises group-theoretic methods that recast disentanglement in terms of the direct product of groups defined on the latent space (Higgins et al., 2018) by equivariant modeling between input and latent vector space rather than 2D planar symmetries. Representative works instantiate the general linear group (GL(n)GL(n)(Zhu et al., 2021; Jung et al., 2024; Winter et al., 2022) or cyclic groups (Yang et al., 2022; Keurti et al., 2023) on latent vector space, and design equivariant VAE objectives that improve disentangled representation quality. However, these models typically retain a fixed standard Gaussian prior, which limits statistical flexibility.
Motivated by these observations, we target two desiderata: 1) preserving partial equivariance between the input and a transformed latent space, and 2) enabling diverse priors. We introduce a learnable, invertible latent-to-latent transformation that preserves partial equivariance between inputs and latent space, and via an exponential-family conversion (EF-conversion), maps the transformed latent vectors to flexible (non-Gaussian) priors. We refer to the module as the Multiple Invertible and Partial-Equivariant (MIPE) transformation, which can be inserted into existing VAEs.
Through quantitative and qualitative experiments on 3D Cars, 3D Shapes, and dSprites, MIPE yields consistent improvements in disentangled representation learning. Our contributions are:
@@
    A symmetric matrix exponential latent-to-latent transformation with invertibility and analyzed partial equivariance properties.
@@
    A training procedure and losses that allow VAEs to learn an unknown latent distribution approximated by an exponential family.
@@
    The MIPE architecture that integrates multiple IPE transformations with EF-conversion and can be plugged into state-of-the-art VAEs.
@@
    Empirical analysis demonstrating the effectiveness of MIPE on standard disentanglement benchmarks.
2 Preliminaries
  • closure: $g_{a},g_{b}\in G\Rightarrow g_{a}\ast g_{b}\in G$.
Group action: Let set $X$, and $(G,\circ)$ be a group, binary operation $\cdot:G\times X\rightarrow X$, then group action $\alpha:\alpha(g,x)=g\cdot x$ following properties:
  • Identity: $e\cdot x=x$, where $e\in G,\,\,x\in X$.
  • Compatibility: $\forall g_{1},g_{2}\in G,\,\,x\in X,\,\,\alpha((g_{1}\circ g_{2}),x)=\alpha(g_{1},\alpha(g_{2},x))$.
Equivariant map: Given $X$ and $Y$ are $G$-set, and group action $\rho:G\times Y\rightarrow Y$. Then a function $f:X\rightarrow Y$ is equivariant if
$f(\alpha(g,x))=\rho(g,f(x)).$ (1)
Partial-Equivariant map: Let subset of $G$ be $G^{\prime}\subset G$, then a function $f:X\rightarrow Y$ is a partial-equivariant if
$f(\alpha(g^{\prime},x))=\rho(g^{\prime},f(x))\,\,\text{where}\,\,g^{\prime}\in G^{\prime}.$ (2)
Homogeneous space: Let a group $G$ act (on the left) on a smooth manifold $X$ via $\cdot:G\times X\to X$. The $G$-space $X$ is called homogeneous if the action is transitive, i.e., for all $x,y\in X$ there exists $g\in G$ with $g\cdot x=y$.
Exponential Family
We introduce the conjugate prior and posterior distribution of the exponential family.
Power Density Function:
$\begin{split}P{{\bm{x}}}({\bm{x}}|\bm{\theta})&=\text{exp}(\bm{\theta}^{\intercal}T({\bm{x}})-A(\bm{\theta})+B({\bm{x}})),\end{split}$ (3)
If observation is a set of independent identically distributed, then Eq. 3 is modified as
$p(\mathbf{X}|\bm{\theta})=\prod_{n=1}^{N}h({\bm{x}}_{n})\text{exp}(\bm{\theta}^{\intercal}\sum_{n=1}^{N}T({\bm{x}}_{n})-A(\bm{\theta})),$ (4)
where observation $\mathbf{X}=\{{\bm{x}}_{1},\cdots{\bm{x}}_{N}\}$ (Bishop, 2006).
Conjugate Prior: Exponential family admits a conjugate prior:
where $B^{\prime}(\cdot)$ is a normalize coefficient and $\nu$ is evidence, and it is expressed by prior natural parameter $\mathcal{\xi}$.
Posterior: Let multiply the prior Eq. 5 by the Eq. 4 to obtain the posterior distribution in the form
$p(\bm{\theta}|\mathbf{X},\mathcal{\xi},\bm{\nu})\propto\text{exp}(\bm{\theta}^{\intercal}(\sum_{n=1}^{N}T({\bm{x}}_{n})+\bm{\nu}\mathcal{\xi})-A(\bm{\theta})),$ (6)
where the evidence $\bm{\nu}$, sufficient statistics $T(\cdot)$ and log-normalizer, $A(\cdot)$ are known functions, samples $\mathbf{X}$ from distribution, and natural parameter of posterior $\bm{\theta}$ and of prior $\mathcal{\xi}$ (Bishop, 2006).
Figure 1: The overall architecture of our proposed MIPET-VAE. The invertible and partial-equivariant function $\psi(\cdot)$ for latent-to-latent (L2L) transformation consists of a symmetric matrix exponential to be 1) invertible and 2) partial-equivariant. Then 3) EF conversion module converges the distribution of unrestricted $\hat{{\bm{z}}}$ to be EF with $\mathcal{L}_{el}$ loss. Also, it applies KL divergence loss ($\mathcal{L}_{kl}$) between the transformed posterior and prior, which are expressed by the power density function of EF. In the last, EF conversion reduces the computational error ($\mathcal{L}_{cali}$) between approximated and true KL divergence.
@@ 3.1 Statistical Approaches for Disentangled Representations
Recently, various studies have focused on unsupervised disentanglement learning. $\beta$-VAE (Higgins et al., 2017) penalizes the Kullback-Leibler divergence (KL divergence) using weighted hyper-parameters. Factor VAE (Kim and Mnih, 2018) and $\beta$-TCVAE (Chen et al., 2018) are trained using total correlation (TC) to encourage independent dimensions in a latent vector, employing a discriminator and decomposed components of the KL divergence term. However, these works are based on the standard Gaussian prior.
@@
Other VAE approaches implement other prior from Gaussian distribution to transformed Gaussian distribution, Gaussian mixture distribution (Dilokthanakul et al., 2016) or von Mises-Fisher distribution (Davidson et al., 2018).  Mathieu et al. (2018) shows that model regularization with a Gaussian prior is inappropriate to convey sufficient dataset information and for disentanglement learning. Rather than using a Gaussian distribution, we employed an exponential family with high flexibility that encompasses a range of different distributions. InteL-VAE (Miao et al., 2022) utilize the invertible function to transform a Gaussian distribution to a complex distribution as Mathieu et al. (2018). Differently, we show a clearer relation of invertibility to disentanglement and improve VAEs to use its unrestricted form of prior. Differently, we consider the recent disentanglement definition based on group theory (Higgins et al., 2018).
3.2 Inductive Bias with Group Theory-Based Approaches
Following the definitions of disentangled representation learning based on group theory, several works have emphasized equivariant and improved disentangled representation learning. Commutative Lie Group VAE (CLG-VAE) (Zhu et al., 2021) proposed a direct mapping of the latent vector into Lie algebra to obtain a group structure (inductive bias) with constraints: commutative and hessian loss. Furthermore, Groupified VAE (Yang et al., 2022) utilizes the Spatial Broadcast Decoder (Watters et al., 2019) to implement an equivariant function to the cyclic group while guaranteeing the commutativity and invertibility of group actions. Topographic VAE (Keller and Welling, 2021) combines Student’s-t distributions with variational inference and enforces rotated latent vectors to be equivariant. On the other hand, we apply an unrestricted prior and posterior for disentanglement learning.
@@ 4 Method
The overview of a VAE equipped with MIPE-transformation is shown in Fig. 1. We aim to explain the three core components of the MIPE-transformation. In Section 4.1, we mathematically prove why the IPE-transformation, which transforms latent vectors, better preserves partial-equivariance when it is an invertible and symmetric matrix. In Section 4.2, we describe EF-conversion, which extends the Gaussian distribution to a diverse exponential family distribution. Finally, in Section 4.3, we illustrate how multiple IPE-transformations and EF-conversions can be integrated into a VAE-based model for disentangled representation.
Notation We denote $g_{I}\in G_{I},g_{L}\in G_{L},\,\text{and}\,g_{T}\in G_{T}$. $M_{n}(\mathbb{R})$ is $n\times n$ real matrix, $E_{M}$ is $\{\mathbf{e}^{M}|M\in M_{n}(\mathbb{R})\}$, $E_{S}$ is a set of $\{\mathbf{e}^{S}|S\in Sym_{n}(\mathbb{R})\}$, and $G_{S}$ is a group $G_{S}:(\mathbf{e}^{S},\ast)$, where $\ast$ is a matrix multiplication (equal to commutative). Further details are in Appendix A.
Assumption Let’s assume that 1) encoder $q_{\phi}(\cdot)$ is a partial-equivariant function, 2) $\mathcal{X},\mathcal{Z}^{\prime},\,\text{and}\,\hat{\mathcal{Z}}^{\prime}$ are homogeneous space of group $G_{I},G_{L},\,\text{and}\,G_{T}$, 3) $G_{L}\subset GL_{n}(\mathbb{R})$, where $GL_{n}(\mathbb{R})$ is the general linear group, and 4) matrices are restricted to a compact subset $K\subset M_{n}(\mathbb{R})$ and probabilities are defined with respect to the normalized uniform probability measure $\mu$ on $K$ (a modeling assumption motivated by finite-precision implementations).
(a) Visualization of homogeneous space (red) as best and worst L2L transformation $\psi(\cdot)$
(b) Venn diagram of groups
Figure 2: The homogeneous space $\mathcal{Z}^{\prime}$ is induced by the encoder $q_{\phi}$, and the cardinality of the $\widehat{\mathcal{Z}}^{\,J}$ depends on the latent-to-latent (L2L) transformation.
4.1 Partial-Equivariant Function and Invertible L2L Transformation
Why Should L2L Transformation be Equivariant?
We transform the latent vector $\bm{z}$} sampled from a Gaussian distribution to the transformed latent vector $\hat{{\bm{z}}}$ to be sampled from a different distribution. However, not all L2L transformations preserve the partial-equivariance of encoder condition. As illustrated in Fig. 2(a), we choose $\psi$ to maximize the cardinality of the homogeneous $G_{T}^{J}$-space induced in latent space:
$\psi^{\star}\;=\;\arg\max_{\psi\in M_{n}(\mathbb{R})}\;\big|\widehat{\mathcal{Z}}^{\,J}(\psi)\big|,\quad$ (7)
where $G_{T}^{J}$ act on the latent space via $\alpha$ and define $\widehat{\mathcal{Z}}^{\,J}(\psi):=\{\alpha(g,\hat{{\bm{z}}})\mid g\in G_{T}^{J}\}$. As a homogeneous $G_{T}^{J}$–space, fixing a base point $\hat{{\bm{z}}}_{0}$ determines all elements $\alpha(g,\hat{{\bm{z}}}_{0})$. If the latent space factors as $\hat{\mathcal{Z}}=\hat{\mathcal{Z}}_{1}\times\cdots\times\hat{\mathcal{Z}}_{k}$ and the action is componentwise, then every $\alpha(g,\hat{{\bm{z}}}_{0})$ with $g\in G_{1}\times\cdots\times G_{k}$ is also disentangled (Higgins et al., 2018).
Equivariance Property with Symmetric Matrix Exponential